# Optimizing Parameters for Static Equilibrium of Discrete Elastic Rods with Active-Set Cholesky

### v1

@@ V-A1 Problems with Naive Substitution of Cholesky Factors
V-A2 Filtered Symmetric Gauss-Seidel
The equivalence between forward (back) substitution and forward (backward) Gauss-Seidel (GS) when applied to lower (upper) triangular matrices allows us to interpret triangular solves as symmetric GS (SGS). Instead of filtering the system matrix $\mathbf{A}$ directly, we apply filtering during SGS to solve a system equivalent to the filtered preconditioning system. Given the elementwise form of forward GS, $\mathbf{y}_{i}=(\mathbf{r}_{i}-\sum_{j<i}\mathbf{L}_{ij}\mathbf{y}_{j})/\mathbf{L}_{ii}$, we can filter this operation with $\mathbf{S}$ while ensuring symmetry using
$\mathbf{y}_{i}=\frac{\mathbf{S}_{ii}\mathbf{r}_{i}-\mathbf{S}_{ii}\sum_{j<i}\mathbf{S}_{jj}\mathbf{L}_{ij}\mathbf{y}_{j}}{\mathbf{L}_{ii}}.$ (32)
@@ V-A2 Filtered Symmetric Gauss-Seidel
$\mathbf{z}_{i}=\frac{\mathbf{S}_{ii}\mathbf{y}_{i}-\mathbf{S}_{ii}\sum_{j>i}\mathbf{S}_{jj}\mathbf{L}_{ji}\mathbf{z}_{j}}{\mathbf{L}_{ii}}.$ (33)
Our filtered SGS preserves the symmetry of operations required for preconditioning of CG/MPRGP [48], and can be rewritten in block form as $(\tilde{\mathbf{D}}+\mathbf{S}\tilde{\mathbf{L}}\mathbf{S}^{T})\mathbf{y}=\mathbf{S}\mathbf{r}$ and $(\tilde{\mathbf{D}}+\mathbf{S}^{T}\tilde{\mathbf{L}}^{T}\mathbf{S})\mathbf{z}=\mathbf{S}\mathbf{y}$, respectively, where $\mathbf{L}=\tilde{\mathbf{D}}+\tilde{\mathbf{L}}$, and $\tilde{\mathbf{D}}$ and $\tilde{\mathbf{L}}$ denote the diagonal and strictly lower parts of $\mathbf{L}$, respectively. Note that while (32) and (33) are equivalent to SGS (except for filtering), our $\mathbf{L}$ arises from Cholesky factorization, in contrast to the strictly lower triangular matrix $\hat{\mathbf{L}}$ from the traditional SGS (where $\mathbf{A}=\hat{\mathbf{D}}+\hat{\mathbf{L}}+\hat{\mathbf{L}}^{T}$, and $\hat{\mathbf{D}}$ is a diagonal part of $\mathbf{A}$). Additionally, while it is typical to initialize $\mathbf{y}=0$ and $\mathbf{z}=0$ in GS-type preconditioning [48], our scheme does not require such initialization because (32) and (33) directly overwrite $\mathbf{y}$ and $\mathbf{z}$.
V-A3 Sqrt-Free Cholesky and Optimization
We prefer sqrt-free Cholesky factorization $\mathbf{A}=\mathbf{L}\mathbf{D}\mathbf{L}^{T}$, where $\mathbf{L}$ is unit lower triangular (for clarity, we redefine $\mathbf{L}$ here), and $\mathbf{D}$ is diagonal. This approach can detect negative pivots (to clamp them for robustness) and avoid sqrt operations for efficiency [49]. The triangular solve proceeds as $\mathbf{L}\mathbf{y}=\mathbf{r}$, $\mathbf{D}\mathbf{w}=\mathbf{y}$, and $\mathbf{L}^{T}\mathbf{z}=\mathbf{w}$. By merging the forward substitution and diagonal scaling (which preserves symmetry) and skipping unnecessary computations (e.g., multiplications with $\mathbf{S}_{ii}$ and $\mathbf{L}_{ii}(=1)$ and scanning rows in $\mathbf{L}$), we have forward and backward operations for $\mathbf{y}$ and $\mathbf{z}$, respectively, as
$\mathbf{y}_{i}$ =$=\begin{cases}(\mathbf{r}_{i}-\sum_{j<i}\mathbf{S}_{jj}\mathbf{L}_{ij}\mathbf{y}_{j})/\mathbf{D}_{ii}&\mathbf{S}_{ii}\neq 0,\\ 0&\mathrm{otherwise},\end{cases}$ (34)
$\mathbf{z}_{i}$ =$=\begin{cases}\mathbf{y}_{i}-\sum_{j>i}\mathbf{S}_{jj}\mathbf{L}_{ji}\mathbf{z}_{j}&\mathbf{S}_{ii}\neq 0,\\ 0&\mathrm{otherwise}.\end{cases}$ (35)
V-B Active-Set Incomplete Cholesky Preconditioning
@@
VI Results and Discussions
We implemented our method in C++20 with double-precision floating-point for scalar values and parallelized forward simulation and parameter optimization with OpenMP, processing each strand concurrently. We executed all the examples on a desktop machine with an Intel Core i7-9700 (8 cores) with 16GB RAM. For forward simulation, we use the exact Hessian with SPD projection for stretching energy and Gauss-Newton Hessian approximation for bending and twisting energies [23]. We perform a single Newton iteration per simulation step [49], executing four simulation steps per frame, except for Figures 1 and 10, which involved 12 simulations steps. We use 60 frames per second. For MPRGP, we use a termination absolute or relative residual of $10^{-10}$. Unless specified otherwise, we use $\mathbf{c}_{\mathrm{st}}=10^{9},\mathbf{c}_{\mathrm{be}}=10^{9},\mathbf{c}_{\mathrm{tw}}=10^{9}$. Material frames are rendered at the centers of the corresponding edges in red and yellow, with lengths matching the rest lengths of the edges.
VI-A Bending With 2D vs. 4D Curvatures
We begin by comparing bending formulations using 2D and 4D curvatures to justify our choice of 4D curvatures, given that our parameter optimization employs only two rest curvatures per vertex. The analysis is conducted on a horizontal strand discretized with 30 vertices.
Figure 2: Comparison of two bending models with a horizontal strand without and with flipped material frames. (a) 2D curvatures. (b) 2D curvatures with flipped frames. (c) 4D curvatures. (d) 4D curvatures with flipped frames. The bending model with 2D and 4D curvatures without edge flips can correctly evaluate bending and generate natural strand behaviors ((a) and (c)). With the flipped frames, the 2D curvature bending model fails to generate bending forces (b), whereas 4D curvature bending model correctly handles the bending, generating the expected result (d).
VI-A1 2D Curvatures with Averaged Material Frames [7]
In Figure 2, we compare a bending model with 4D curvatures [6] against a bending model with 2D curvatures [7] which averages material frames to reduce curvature dimensions from 4D to 2D. We experiment with the horizontal strand without and with its material frames flipped, and use $\mathbf{c}_{\mathrm{be}}=10^{10}$.
While both models without edge flips yield equivalent results, the 2D curvature model [7] fails to generate bending forces when material frames are flipped because the averaged material frames can be non-unit vectors (in this example, averaged frames are zero vectors) [9, 12, 41]. By contrast, the 4D curvature model [6] correctly evaluates the bending even with the flipped frames, producing results consistent with those obtained without edge flips.
Figure 3: Comparison of two bending models with a horizontal strand. 2D curvatures with slerp and $\mathbf{c}_{\mathrm{be}}=10^{8}$ (a) and $\mathbf{c}_{\mathrm{be}}=10^{9}$ (b). 4D curvatures with $\mathbf{c}_{\mathrm{be}}=10^{8}$ (c) and $\mathbf{c}_{\mathrm{be}}=10^{9}$ (d). While spherical interpolation enables correct evaluation of bending even with the flipped material frames (b), approximated gradients can lead to stability problems (a). In contrast, the bending model with 4D curvatures generates natural strand behaviors ((c) and (d)).
VI-A2 2D Curvatures with Spherically Interpolated Material Frames [41]
Figure 3 contrasts the bending model with 4D curvatures [6] against another model [41] with 2D curvatures, which spherically interpolates (i.e., uses slerp) material frames to maintain unit-size frames while reducing the curvature dimensions. We use flipped frames and $\mathbf{c}_{\mathrm{be}}=10^{8},10^{9}$.
While the slerped material frames, which retain unit length, enable correct bending evaluation even with the flipped frames, differentiating the slerped frames to compute bending forces presents challenges. Consequently, their bending model [41] resorts to approximated gradients, which proved insufficiently accurate in our experiments, causing stability problems with the Gauss-Newton Hessian approximation, whereas the bending model of [6] is stable under the same settings.
@@ VI Results and Discussions
Figure 4: Coil-like strand test. Rest shape optimization achieves static equilibrium unlike naive initialization (a). 4D rest curvatures with unbanded/banded systems generate the identical results (b) while the reduced 2D rest curvatures also produce equivalent results given the redundant force spaces (c).
VI-B Banded System with Reduced Rest Curvatures
To assess the performance improvement achieved through banded systems and reduced rest curvatures, we experiment with a coil-like strand discretized with 2k vertices, as shown in Figure 4 (we include naive initialization for reference). For a fair comparison with previous work [5], we optimize only the rest shape parameters and exclude box constraints to eliminate the influence of active sets. We compare the following:
@@ VI-B Banded System with Reduced Rest Curvatures
Due to the redundant force spaces with rest curvatures, the virtually reduced 2D rest curvatures produce results comparable to those obtained with 4D rest curvatures. The computation times were 2.1ms and 5.0ms with 2D and 4D rest curvatures, respectively. As the DOF count is $84N-$ 8 and $126N-$ 12, with non-zero counts per row of $22.0$ and $32.0$ on average, the total number of non-zeros is approximately $88N(=4N\times 22)$ and $192N(=6N\times 32)$ for 2D and 4D rest curvatures, respectively. Thus, the performance ratio $0.42=(2.1/5.0)$ is in good agreement with the ratio of the total number of non-zeros $0.46\approx(88N/192N)$.
@@
Figure 5: Comparison with a horizontal strand. The penalty method fails to precisely enforce box constraints, causing significant rest curvature changes and allowing the strand’s tail to end up on the left side after perturbing the root end (b). By contrast, our active-set method (c) strictly enforces the box constraints, keeping the tail on the right side and better preserving the original shape than the naive initialization (a).
VI-C Penalty vs. Active-Set Methods
To demonstrate the efficacy of our active-set-based BCQP solver, we compare it to a penalty-based one, using a horizontal strand discretized with 30 vertices, as shown in Figure 5 (naive initialization included for reference). In this comparison, we optimize rest shape parameters only and use $\mu=0.4$. For the penalty method, we use a Cholesky solver [5]. As we perform only one Newton iteration to solve each BCQP, if we apply ALM (with Lagrange multipliers initialized to zero) to the BCQP, it becomes equivalent to the penalty method [15].
With the penalty method, it is difficult to accurately control rest curvature changes, allowing rest curvatures to violate the box constraints. Consequently, significant rest shape changes caused the tail end of the strand to flip to the left side after the root was perturbed. In contrast, our active-set method precisely constrains curvature changes retaining the tail end of the strand on the right side. While we may need to compromise static equilibrium, our method better preserves the original shape compared to naive initialization.
@@
VI-D Simultaneous Optimization of Rest Shape and Material Stiffness Parameters Under Hard Zero Net Force Constraints
@@
    Rest shape and material stiffness with penalty: simultaneous optimization of rest shape and material stiffness parameters with enforcement of $\mathbf{c}(\mathbf{p})=0$ via quadratic penalty [15] (which is equivalent to ours with $\bm{\lambda}=0$);
@@
Figure 6: Evaluation with a vertical strand. The first black edge indicates the invalid material stiffness parameter for stretching. The white and green edges represent lower and higher stiffness parameters, respectively. (a) Rest shape only. (b) Rest shape and material stiffness with penalty. (c) Rest shape and material stiffness with ALM (ours). With the rest shape only optimization, material stiffness parameters are unchanged (edge stay white), failing to achieve static equilibrium with rest shape parameters within their box constraints (a). Optimizing the material stiffness parameters additionally stiffens the edges and thus reduces the necessary rest shape changes. Treating the zero net force constraint as a hard constraint enables more significant stiffness changes to achieve static equilibrium (c), compared to using a soft constraint (b).
VI-D1 Vertical Strand
@@
The rest shape only optimization uses 8 Newton iterations, 5.6 MPRGP iterations on average, and takes 8.1 ms for the entire optimization process, while the rest shape and material stiffness optimization with penalty uses 7 Newton iterations, 73.9 MPRGP iterations, and takes 19.5 ms for the optimization. The increased cost is primarily due to the larger system size ($147N-$ 14 for rest shape and material stiffness vs. $84N-$ 8 for rest shape only) with more non-zeros and increased MPRGP iterations (due to the increased complexity of the systems). Our method uses 25 Newton iterations, 66.0 MPRGP iterations, and takes 58.3ms for the optimization. The increased cost compared to the approach with penalty is due to the updates of Lagrange multipliers, which modifies the optimization problem. Consequently, our method requires around 3×3$\times$3 × more Newton iterations and computational time. Although our method is approximately 7×7$\times$7 × more costly than the rest shape only optimization [5], these methods are performed only once per scene as a preprocess (i.e., no additional cost during forward simulation). Thus, we believe that our approach, which guarantees static equilibrium with minimal stiffness changes while eliminating tedious manual stiffness tuning, is a practical and attractive alternative.
@@
VI-D2 Horizontal Strand
We experiment with a horizontal strand discretized with 30 vertices (Figure 7) using $\mu=0.2$. Similar to the vertical strand case, optimizing only the rest shape parameters fails to achieve static equilibrium (despite the relatively stiff strand with $\mathbf{c}_{\mathrm{be}}=10^{9}$) due to tightly bounded rest curvature changes. While simultaneous optimization of the rest shape and material stiffness aims to achieve zero net forces, it still fails when enforcing $\mathbf{c}(\mathbf{p})=0$ as a soft constraint. By contrast, our method successfully attains static equilibrium. The computational costs follow a trend similar to that observed with the vertical strand. The optimization takes 7.5 ms for the rest shape only, 17.9 ms for the rest shape and material stiffness with penalty, and 47.7 ms for our method.
VI-E Box-Constrained Quadratic Program Solver Comparisons

### v2

@@ VI-A1 Problems with Naive Application of Cholesky Solver
VI-A2 Our Preconditioning
Considering the equivalence between forward/back substitution and forward/backward GS applied to lower/upper triangular matrices, we can interpret triangular solves as SGS. Thus, instead of directly filtering the system matrix $\mathbf{A}$ with $\mathbf{S}$, we perform the filtering in SGS to solve a system equivalent to the filtered preconditioning system. Given the elementwise form of forward GS, $\mathbf{y}_{i}=(\mathbf{r}_{i}-\sum_{j<i}\mathbf{L}_{ij}\mathbf{y}_{j})/\mathbf{L}_{ii}$, we can filter this operation with $\mathbf{S}$ while ensuring symmetry by
$\mathbf{y}_{i}=(\mathbf{S}_{ii}\mathbf{r}_{i}-\mathbf{S}_{ii}\sum_{j<i}\mathbf{S}_{jj}\mathbf{L}_{ij}\mathbf{y}_{j})/\mathbf{L}_{ii}.$ (20)
@@ VI-A2 Our Preconditioning
$\mathbf{z}_{i}=(\mathbf{S}_{ii}\mathbf{y}_{i}-\mathbf{S}_{ii}\sum_{j>i}\mathbf{S}_{jj}\mathbf{L}_{ji}\mathbf{z}_{j})/\mathbf{L}_{ii}.$ (21)
Note that it is unnecessary to filter the denominator $\mathbf{L}_{ii}$ in (20) and (21) because $\mathbf{y}$ and $\mathbf{z}$ are properly filtered in the numerator. Our filtered SGS preserves the symmetry of operations required for preconditioning of CG/MPRGP [52], and can be rewritten in the block form as $(\tilde{\mathbf{D}}+\mathbf{S}\tilde{\mathbf{L}}\mathbf{S}^{T})\mathbf{y}=\mathbf{S}\mathbf{r}$ and $(\tilde{\mathbf{D}}+\mathbf{S}^{T}\tilde{\mathbf{L}}^{T}\mathbf{S})\mathbf{z}=\mathbf{S}\mathbf{y}$, respectively, where $\mathbf{L}=\tilde{\mathbf{D}}+\tilde{\mathbf{L}}$, and $\tilde{\mathbf{D}}$ and $\tilde{\mathbf{L}}$ denote the diagonal and strictly lower parts of $\mathbf{L}$, respectively. Note that while (20) and (21) are equivalent to SGS (except for filtering), our $\mathbf{L}$ arises from Cholesky factorization, in contrast to the strictly lower triangular matrix $\hat{\mathbf{L}}$ from the traditional SGS (where $\mathbf{A}=\hat{\mathbf{D}}+\hat{\mathbf{L}}+\hat{\mathbf{L}}^{T}$, and $\hat{\mathbf{D}}$ is the diagonal part of $\mathbf{A}$). Additionally, while it is typical to initialize $\mathbf{y}=0$ and $\mathbf{z}=0$ in GS-type preconditioning [52], our scheme does not require such initialization because (20) and (21) directly overwrite $\mathbf{y}$ and $\mathbf{z}$.
In practice, we prefer sqrt-free Cholesky factorization $\mathbf{A}=\mathbf{L}\mathbf{D}\mathbf{L}^{T}$, where $\mathbf{L}$ is unit lower triangular (for clarity, we redefine $\mathbf{L}$ here), and $\mathbf{D}$ is diagonal. This approach can detect negative pivots (to clamp them for robustness) and avoid sqrt operations for efficiency [47]. The triangular solve proceeds as $\mathbf{L}\mathbf{w}=\mathbf{r}$, $\mathbf{D}\mathbf{y}=\mathbf{w}$, and $\mathbf{L}^{T}\mathbf{z}=\mathbf{y}$. By merging the forward substitution and diagonal scaling (which preserves symmetry) and skipping unnecessary computations (e.g., multiplications with $\mathbf{S}_{ii}$ and $\mathbf{L}_{ii}(=1)$, and scanning rows in $\mathbf{L}$), we have forward and backward operations for $\mathbf{y}$ and $\mathbf{z}$, respectively, as
$\mathbf{y}_{i}$ $=\begin{cases}(\mathbf{r}_{i}-\sum_{j<i}\mathbf{S}_{jj}\mathbf{L}_{ij}\mathbf{y}_{j})/\mathbf{D}_{ii}&\mathbf{S}_{ii}\neq 0,\\ 0&\mathrm{otherwise},\end{cases}$ (22)
$\mathbf{z}_{i}$ $=\begin{cases}\mathbf{y}_{i}-\sum_{j>i}\mathbf{S}_{jj}\mathbf{L}_{ji}\mathbf{z}_{j}&\mathbf{S}_{ii}\neq 0,\\ 0&\mathrm{otherwise}.\end{cases}$ (23)
VI-B Active-Set Incomplete Cholesky Preconditioning
@@
VII Results and Discussions
We implemented our method in C++20 with double-precision floating-point for scalar values and parallelized forward simulation and parameter optimization with OpenMP, processing each strand concurrently. We executed all the examples on a desktop machine with an Intel Core i7-9700 (8 cores) with 16GB RAM. For forward simulation, we use the exact Hessian with SPD projection for stretching objective and Gauss-Newton Hessian approximation for bending and twisting objectives [23]. We perform a single Newton iteration per simulation step [47], executing four simulation steps per frame, except for Figures 1 and 8, which used 12 simulation steps per frame. We use 60 frames per second. For MPRGP, we use a termination absolute or relative residual of $10^{-10}$. Unless specified otherwise, we use $\mathbf{c}_{\mathrm{st}}=10^{9},\mathbf{c}_{\mathrm{be}}=10^{9},\mathbf{c}_{\mathrm{tw}}=10^{9}$. Material frames are rendered at the centers of the corresponding edges in red and yellow, with lengths matching the rest lengths of the edges.
@@ VII Results and Discussions
Figure 2: Coil-like strand test. Rest shape optimization achieves static equilibrium unlike naive initialization (a). 4D rest curvatures with unbanded/banded systems generate identical results (b), while the reduced 2D rest curvatures also produce equivalent results given the redundant force spaces (c).
VII-A Banded System with Reduced Rest Curvatures
To assess the performance improvement achieved through banded systems and reduced rest curvatures, we experiment with a coil-like strand discretized with 2k vertices, as shown in Figure 2 (for reference, we include naive initialization, which computes rest shape parameters with the generalized positions of the input strand). For a fair comparison with previous work [5], we optimize only the rest shape parameters and exclude box constraints to eliminate the influence of active sets. We compare the following:
@@ VII-A Banded System with Reduced Rest Curvatures
Due to the redundant force spaces with rest curvatures, the virtually reduced 2D rest curvatures produce results comparable to those obtained with 4D rest curvatures. The computation times were 2.1 ms and 5.0 ms with 2D and 4D rest curvatures, respectively. As the DOF count is 4N84N-8 and 6N126N-12, with non-zero counts per row of $22.0$ and $32.0$ on average, the total number of non-zeros is approximately $88N(=4N\times 22)$ and $192N(=6N\times 32)$ for 2D and 4D rest curvatures, respectively. Thus, the performance ratio $0.42=(2.1/5.0)$ is in good agreement with the ratio of the total number of non-zeros $0.46\approx(88N/192N)$.
@@
Figure 3: We vertically translate the root position (initially the leftmost endpoint) of a horizontal strand up and down. The leftmost column shows the initialized states for each method and the remaining columns from left to right show the results of forward simulation. The penalty method fails to precisely enforce box constraints, causing significant rest curvature changes and allowing the strand’s tail to end up on the left side after the motion stops (b). By contrast, our active-set method (c) strictly enforces the box constraints, keeping the tail on the right side and better preserving the original shape than the naive initialization (a).
VII-B Penalty vs. Active-Set Methods
To demonstrate the efficacy of our active-set-based BCQP solver, we compare it to a penalty-based one, using a horizontal strand discretized with 30 vertices, as shown in Figure 3 (naive initialization included for reference). In this comparison, we optimize rest shape parameters only and use $\mu=0.4$ to prevent significant rest curvature changes and thus to keep the tail end of the strand on the right. For the penalty method, we use a Cholesky solver [5]. As we perform only one Newton iteration to solve each BCQP, if we apply ALM (with Lagrange multipliers initialized to zero) to the BCQP, it becomes equivalent to the penalty method [15].
The penalty method failed to strictly enforce the rest curvature changes within their corresponding box constraints. Consequently, significant rest shape changes caused the tail end of the strand to flip to the left side after the root was perturbed. By contrast, our active-set method precisely constrains curvature changes via their box constraints, keeping the tail end of the strand on the right side. While we may need to compromise static equilibrium, our method better preserves the original shape compared to naive initialization.
@@
VII-C Simultaneous Optimization of Rest Shape and Material Stiffness Parameters Under Hard Zero Net Force Constraints
@@
    Rest shape and material stiffness with penalty: simultaneous optimization of rest shape and material stiffness parameters with enforcement of $\mathbf{c}(\mathbf{p})=0$ via quadratic penalty [15] (which is equivalent to ours with $\boldsymbol{\lambda}=0$);
@@
Figure 4: Evaluation with a vertical strand. The first edge (black) is fixed so its stretching stiffness parameter is undefined. The white/green edges represent lower/higher stiffness parameters. (a) Rest shape only. (b) Rest shape and material stiffness with penalty. (c) Rest shape and material stiffness with ALM (ours). With the rest shape only optimization, material stiffness parameters are unchanged (edges stay white), failing to achieve static equilibrium with rest shape parameters within their box constraints (a). Optimizing the material stiffness parameters additionally stiffens the edges and thus reduces the necessary rest shape changes. Treating the zero net force constraint as a hard constraint enables more significant stiffness changes to achieve static equilibrium (c), compared to using a soft one (b).
VII-C1 Vertical Strand
@@
The rest shape only optimization takes 8 Newton iterations, 5.6 MPRGP iterations on average, and 8.1 ms for the entire optimization process, while the rest shape and material stiffness optimization with penalty takes 7 Newton iterations, 73.9 MPRGP iterations, and 19.5 ms for the optimization. The increased cost is primarily due to the larger system size (7N147N-14 for rest shape and material stiffness vs. 4N84N-8 for rest shape only) with more non-zeros and increased MPRGP iterations (due to the increased complexity of the systems). Our method takes 25 Newton iterations, 66.0 MPRGP iterations, and 58.3 ms for the optimization. The increased cost compared to the approach with penalty is due to the updates of Lagrange multipliers, which modify the optimization problem. Consequently, our method requires around $3\times$ more Newton iterations and computational time. Although our method is approximately $7\times$ more costly than the rest shape only optimization [5], these methods are performed only once per scene as a preprocess (i.e., no additional cost during forward simulation). Thus, we believe that our approach, which guarantees static equilibrium with minimal stiffness changes while eliminating tedious manual stiffness tuning, is a practical and attractive alternative.
@@
VII-C2 Horizontal Strand
We experiment with a horizontal strand discretized with 30 vertices (Figure 5) using $\mu=0.2$. Similar to the vertical strand case, optimizing only the rest shape parameters fails to achieve static equilibrium (despite the relatively stiff strand with $\mathbf{c}_{\mathrm{be}}=10^{9}$) due to tightly bounded rest curvature changes. While simultaneous optimization of the rest shape and material stiffness aims to achieve zero net forces, it still fails when enforcing $\mathbf{c}(\mathbf{p})=0$ as a soft constraint using the penalty method. By contrast, our method successfully attains static equilibrium. The computational costs follow a trend similar to that observed with the vertical strand. The optimization takes 7.5 ms for the rest shape only, 17.9 ms for the rest shape and material stiffness with penalty, and 47.7 ms for our method.
VII-D Box-Constrained Quadratic Program Solver Comparisons